# Aditya Tanna, Pratinav Seth, Mohamed Bouadi Utsav Avaiya, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v2

@@ 6.2.1  Overall Leaderboards
Episodic meta-learning substantially improves both accuracy and rank stability relative to zero-shot inference. OrionMSP attains the lowest aggregate rank of 2.26, achieving (ACC=0.8401/F1=0.8310) on TALENT, 0.8548/0.8516 on OpenML-CC18, and (0.8735/0.8672) on TabZilla. TabPFN ranks 2.42 overall with accuracies of (0.8517/0.8414) on TALENT, (0.8842/0.8784) on OpenML-CC18, and (0.8663/0.8603) on TabZilla. TabDPT (rank 3.95) maintains consistent performance near (0.83/0.82). By contrast, architectures such as TabICL and OrionBiX exhibit instability on TabZilla, where TabICL’s F1 drops to 0.6845. These results indicate that meta-learning offers the most balanced compromise between generalization and task-specific adaptation.
Supervised Fine-Tuning (SFT).
@@
Full parameter fine-tuning yields divergent outcomes across architectures. TabPFN benefits most, achieving the best overall rank (1.97) and maintaining high accuracy across all benchmarks (ACC = 0.8459–0.8697, F1 = 0.8350–0.8617). TabDPT ranks 2.79 overall and performs particularly well on TabZilla (rank 1.81), with (ACC = 0.8337, F1 = 0.8260). OrionMSP remains competitive (rank 2.88) with (ACC $\approx$ 0.79, F1 $\approx$ 0.76), whereas TabICL and OrionBiX suffer severe degradation—TabICL’s accuracy on TabZilla falls from 0.873 to 0.567 and F1 from 0.870 to 0.473 —indicating overfitting and loss of pretrained priors. These findings suggest that SFT is advantageous for Bayesian or probabilistic architectures such as TabPFN.
Parameter-Efficient Fine-Tuning (PEFT).
@@
PEFT achieves near-SFT accuracy with significantly reduced computational overhead. In the meta-learning configuration, OrionMSP attains the best overall rank (2.21), recording (ACC = 0.7879/F1 = 0.7728) on TALENT, (0.8566/0.8453) on OpenML-CC18, and (0.8594/0.8581) on TabZilla. TabDPT follows closely (rank 2.28) with (0.8002/0.7910) on TALENT and (0.8600/0.8539) on OpenML-CC18. Under the supervised PEFT variant, TabDPT achieves the overall best rank (1.91) with (ACC $\approx$ 0.85 and F1 $\approx$ 0.84), while OrionMSP (rank 2.01) attains (0.821 / 0.807) on TabZilla. These results show that PEFT retains roughly 95% of full-fine-tuning accuracy while substantially lowering resource demands.
Discussion.
Across all adaptation regimes, TFMs consistently outperform traditional baselines by 2–4 percentage points in accuracy and approximately four rank positions.
TabPFN proves most resilient, achieving the top overall rank under SFT, whereas OrionMSP demonstrates superior cross-dataset generalization in meta-learning. TabDPT delivers near-optimal performance under PEFT, offering the most favorable efficiency–accuracy balance.
@@ Discussion.
Conversely, models such as TabICL and OrionBiX are highly sensitive to full parameter updates and prone to overfitting. Collectively, these observations establish that careful alignment between model architecture and adaptation strategy is critical for maximizing the performance of tabular foundation models.
@@
  • Overall leaderboard: Considering scores across all benchmark suites, the strongest overall performers are TabPFN and OrionMSP. TabPFN achieves the highest overall results, with mean accuracies between 0.85–0.88 and weighted F1 scores of 0.84–0.88, demonstrating exceptional cross-benchmark consistency. OrionMSP follows closely, reaching ACC $\approx$ 0.84–0.88 and F1 $\approx$ 0.83–0.88, and exhibits particularly strong generalization on OpenML-CC18 and TabZilla. Collectively, these Tabular Foundation Models (TFMs) surpass classical machine learning baselines by approximately 2–4 %, underscoring the advantage of pretrained tabular representations for cross-dataset generalization.
  • Zero-shot performance: TabPFN and OrionMSP are the strongest zero-shot models, achieving $\text{ACC}\approx 0.85$$0.88$ and $\text{F1}\approx 0.84$$0.88$. OrionMSP attains peak accuracy on OpenML-CC18 and TabZilla, while TabPFN leads on TALENT and remains stable across all suites.
  • Meta-learning benefits: Meta-learning fine-tuning consistently improves both accuracy and stability. OrionMSP achieves the best meta-learning results overall ($\text{ACC}\approx 0.84$$0.87$, $\text{F1}\approx 0.83$$0.87$), with TabPFN performing comparably and showing particularly strong results on OpenML-CC18.
@@ 6.2.2  Scalability Analysis
We analyze how model performance varies with dataset size, feature dimensionality, and class imbalance to inform model selection under different data regimes. This scalability analysis highlights which models and fine-tuning strategies perform best across varying data constraints.
@@
Models / Strategy Small (<1K) Medium (1K-10K) Large (>10K)
ACC F1 ACC F1 ACC F1
@@
XGBoost 0.8168 0.7964 0.8363 0.8314 0.8969 0.8920
CatBoost 0.8124 0.7935 0.8340 0.8264 0.8797 0.8733
Random Forest 0.7988 0.8187 0.8285 0.8221 0.8694 0.8628
LightGBM 0.8143 0.7789 0.8314 0.8226 0.8827 0.8764
TabICL 0.8301 0.8338 0.8486 0.8398 0.8802 0.8743
OrionBiX 0.8330 0.8150 0.8348 0.8260 0.8729 0.8670
OrionMSP 0.8232 0.8194 0.8494 0.8402 0.8843 0.8768
TabPFN 0.8325 0.8131 0.8557 0.8462 0.8783 0.8713
Mitra 0.4334 0.3236 0.3600 0.2553 0.3837 0.2754
ContextTab 0.7578 0.7363 0.6210 0.5566 0.6388 0.5638
TabDPT 0.8333 0.8271 0.8424 0.8339 0.8831 0.8765
@@
TabICL 0.7912 0.7806 0.8342 0.8256 0.8431 0.8360
OrionBiX 0.8149 0.8123 0.8283 0.8204 0.8410 0.8306
OrionMSP 0.8211 0.8151 0.8413 0.8330 0.8787 0.8721
TabPFN 0.8336 0.8256 0.8638 0.8548 0.8748 0.8668
Mitra 0.6417 0.5988 0.6083 0.5516 0.6585 0.6132
TabDPT 0.8224 0.8188 0.8289 0.8212 0.8682 0.8613
@@
TabICL 0.6746 0.6225 0.7553 0.7187 0.6693 0.6212
OrionBiX 0.7158 0.6792 0.7718 0.7386 0.6697 0.6272
OrionMSP 0.8135 0.7920 0.8036 0.7798 0.7231 0.6828
TabPFN 0.8216 0.8109 0.8580 0.8485 0.8572 0.8459
Mitra 0.5625 0.4408 0.5133 0.4116 0.5720 0.4501
TabDPT 0.8227 0.8153 0.8377 0.8284 0.8120 0.8002
@@
TabICL 0.7518 0.7397 0.7251 0.7114 0.6962 0.6799
OrionBiX 0.8093 0.8042 0.7995 0.7939 0.7848 0.7665
OrionMSP 0.8275 0.8213 0.8134 0.8019 0.8013 0.7817
Mitra 0.6510 0.6012 0.6089 0.5599 0.6304 0.5795
TabDPT 0.8191 0.8157 0.8121 0.8049 0.8407 0.8299
@@
TabICL 0.7957 0.7710 0.6424 0.5686 0.6084 0.5020
OrionBiX 0.7576 0.7102 0.6585 0.5892 0.6327 0.5381
OrionMSP 0.8298 0.8132 0.7190 0.6597 0.6931 0.6152
Mitra 0.5244 0.4133 0.5008 0.4013 0.5459 0.4476
TabDPT 0.8269 0.8182 0.8332 0.8245 0.7686 0.7418
@@
  • On small datasets (<<1K samples), TabDPT achieves the highest zero-shot accuracy (ACC=0.8333, F1=0.8271), marginally ahead of TabPFN under meta-learning (0.8336/0.8256). OrionMSP (PEFT–Meta, 0.8275/0.8213) follows closely, confirming that transformer backbones still generalize well under constrained data. Fine-tuning TabICL or OrionBiX on small datasets causes sharp degradation (ACC drops >>15%), revealing overfitting risk.
@@
  • On large-scale datasets (>>10K), tree-based ensembles remain strong—XGBoost reaches (0.8969/0.8920)—but transformer TFMs scale competitively, with OrionMSP zero-shot (0.8843/0.8768) and TabDPT zero-shot (0.8831/0.8765) ranking second and third. Across all sizes, Mitra and ContextTab consistently underperform (ACC <0.65<0.65), reflecting limited scalability.
@@
Models / Strategy Narrow (<10) Medium (10-100) Wide (>100)
ACC F1 ACC F1 ACC F1
@@
XGBoost 0.8222 0.8159 0.8482 0.8410 0.9140 0.9039
CatBoost 0.8145 0.8067 0.8441 0.8344 0.9157 0.9084
Random Forest 0.8005 0.7044 0.8410 0.8235 0.9034 0.8936
LightGBM 0.8128 0.7907 0.8458 0.8326 0.8999 0.8908
TabICL 0.8208 0.8119 0.8627 0.8549 0.9101 0.8936
OrionBiX 0.8112 0.8043 0.8510 0.8417 0.8859 0.8849
OrionMSP 0.8394 0.8314 0.8572 0.8478 0.8860 0.8837
TabPFN 0.8187 0.8092 0.8676 0.8589 0.9129 0.9111
Mitra 0.3737 0.2683 0.3886 0.2781 0.2521 0.1497
ContextTab 0.6391 0.5719 0.6480 0.5843 0.6017 0.5651
TabDPT 0.8262 0.8189 0.8566 0.8483 0.8845 0.8820
@@
TabICL 0.7867 0.7779 0.8383 0.8338 0.8319 0.8667
OrionBiX 0.7956 0.7900 0.8316 0.8278 0.8429 0.8640
OrionMSP 0.8253 0.8190 0.8514 0.8457 0.8564 0.8737
TabPFN 0.8264 0.8171 0.8683 0.8579 0.9172 0.9533
Mitra 0.6646 0.6247 0.6004 0.5763 0.4090 0.3224
TabDPT 0.8129 0.8071 0.8405 0.8350 0.8400 0.8497
@@
TabICL 0.6609 0.6195 0.7507 0.7101 0.7757 0.7269
OrionBiX 0.6922 0.6576 0.7643 0.7294 0.7560 0.7046
OrionMSP 0.7540 0.7297 0.8025 0.7750 0.8148 0.7808
TabPFN 0.8258 0.8162 0.8578 0.8469 0.9346 0.9335
Mitra 0.5619 0.4478 0.5349 0.4268 0.3307 0.2436
TabDPT 0.8078 0.7948 0.8378 0.8294 0.8623 0.8587
@@
TabICL 0.6375 0.6375 0.7601 0.7444 0.7714 0.7622
OrionBiX 0.7584 0.7584 0.8147 0.8057 0.8251 0.8243
OrionMSP 0.7836 0.7836 0.8254 0.8107 0.8293 0.8256
Mitra 0.6471 0.6471 0.6177 0.5650 0.4129 0.3338
TabDPT 0.7979 0.7979 0.8299 0.8209 0.8416 0.8376
@@
TabICL 0.6196 0.5287 0.6596 0.5902 0.8760 0.8502
OrionBiX 0.6278 0.5444 0.6851 0.6177 0.6737 0.6166
OrionMSP 0.6795 0.6113 0.7477 0.6939 0.8042 0.7663
Mitra 0.5518 0.4481 0.5106 0.4108 0.2831 0.1874
TabDPT 0.7821 0.7631 0.8305 0.8201 0.8925 0.8901
@@
  • For narrow datasets (<<10 features), performance converges across models: OrionMSP zero-shot (0.8394/0.8314) slightly leads, followed by TabDPT and TabPFN, while classical baselines remain competitive.
  • On medium-width data (10–100 features), TabPFN dominates under both zero-shot (0.8676/0.8589) and meta-learning (0.8683/0.8579), outperforming all baselines and transformer variants.
  • The advantage of TFMs becomes most pronounced on wide feature spaces (>>100 features): TabPFN achieves exceptional results under SFT (ACC=0.9346, F1=0.9335) and meta-learning (0.9172/0.9533), far exceeding other models. TabICL and OrionBiX degrade when fully fine-tuned, confirming that dense feature spaces accentuate overfitting. Traditional models (CatBoost, XGBoost) remain competitive but trail by 2–3 %.
@@
XGBoost 7.00 0.8175 0.8110 6.23 0.8859 0.8785
CatBoost 7.15 0.8076 0.8020 5.65 0.8785 0.8665
Random Forest 7.92 0.7983 0.7955 6.77 0.8741 0.8646
LightGBM 7.32 0.8071 0.7977 6.19 0.8775 0.8633
TabICL 4.72 0.8279 0.8233 5.08 0.8806 0.8698
OrionBiX 5.65 0.8096 0.8040 5.04 0.8787 0.8683
OrionMSP 4.22 0.8265 0.8202 3.38 0.8840 0.8731
TabPFN 3.85 0.8367 0.8309 5.37 0.8808 0.8697
Mitra 12.26 0.2763 0.1540 11.24 0.4794 0.3858
ContextTab 9.66 0.5079 0.4487 9.72 0.7850 0.7192
TabDPT 5.16 0.8233 0.8189 5.65 0.8798 0.8690
@@
Discussion.
Scalability analysis underscores that TabPFN, OrionMSP, and TabDPT each occupy distinct strengths. TabPFN excels on small, medium, and high-dimensional data where Bayesian priors and strong supervision prevail; OrionMSP scales best with large datasets and imbalanced distributions due to meta-learning robustness; and TabDPT sustains near-top performance across all settings with minimal computational overhead. Classical baselines remain reliable but consistently trail the best TFMs by 2–4% in accuracy.
@@ Discussion.
  • Dataset size: Small datasets favor TabDPT (zero-shot) and TabPFN (Meta Learning); medium datasets favor TabPFN (meta & SFT); large datasets see XGBoost slightly leading, with OrionMSP and TabDPT close behind.
@@
  • Class imbalance: Meta-learning fine-tuning yields the best robustness under imbalance, with OrionMSP (Rank-$2.16$, ACC = $0.8735$, F1 = $0.8636$) outperforming other TFMs; TabPFN (rank $2.50$) remains second-best.

### v3

@@ 6.2.1  Overall Leaderboards
Meta-Learning Fine-Tuning : Episodic meta-learning substantially improves both accuracy and rank stability relative to zero-shot inference. OrionMSP attains the lowest aggregate rank of 2.26, achieving (ACC=0.8401/F1=0.8310) on TALENT, 0.8548/0.8516 on OpenML-CC18, and (0.8735/0.8672) on TabZilla. TabPFN ranks 2.42 overall with accuracies of (0.8517/0.8414) on TALENT, (0.8842/0.8784) on OpenML-CC18, and (0.8663/0.8603) on TabZilla. TabDPT (rank 3.95) maintains consistent performance near (0.83/0.82). By contrast, architectures such as TabICL and OrionBiX exhibit instability on TabZilla, where TabICL’s F1 drops to 0.6845. These results indicate that meta-learning offers the most balanced compromise between generalization and task-specific adaptation.
@@
Supervised Fine-Tuning (SFT) : Full parameter fine-tuning yields divergent outcomes across architectures. TabPFN benefits most, achieving the best overall rank (1.97) and maintaining high accuracy across all benchmarks (ACC = 0.8459–0.8697, F1 = 0.8350–0.8617). TabDPT ranks 2.79 overall and performs particularly well on TabZilla (rank 1.81), with (ACC = 0.8337, F1 = 0.8260). OrionMSP remains competitive (rank 2.88) with (ACC $\approx$ 0.79, F1 $\approx$ 0.76), whereas TabICL and OrionBiX suffer severe degradation—TabICL’s accuracy on TabZilla falls from 0.873 to 0.567 and F1 from 0.870 to 0.473 —indicating overfitting and loss of pretrained priors. These findings suggest that SFT is advantageous for Bayesian or probabilistic architectures such as TabPFN.
@@
Parameter-Efficient Fine-Tuning (PEFT) : PEFT achieves near-SFT accuracy with significantly reduced computational overhead. In the meta-learning configuration, OrionMSP attains the best overall rank (2.21), recording (ACC = 0.7879/F1 = 0.7728) on TALENT, (0.8566/0.8453) on OpenML-CC18, and (0.8594/0.8581) on TabZilla. TabDPT follows closely (rank 2.28) with (0.8002/0.7910) on TALENT and (0.8600/0.8539) on OpenML-CC18. Under the supervised PEFT variant, TabDPT achieves the overall best rank (1.91) with (ACC $\approx$ 0.85 and F1 $\approx$ 0.84), while OrionMSP (rank 2.01) attains (0.821 / 0.807) on TabZilla. These results show that PEFT retains roughly 95% of full-fine-tuning accuracy while substantially lowering resource demands.
@@
Discussion : Across all adaptation regimes, TFMs consistently outperform traditional baselines by 2 % points in accuracy and approximately four rank positions. TabPFN proves most resilient, achieving the top overall rank under SFT, whereas OrionMSP demonstrates superior cross-dataset generalization in meta-learning. TabDPT delivers near-optimal performance under PEFT, offering the most favorable efficiency–accuracy balance. Conversely, models such as TabICL and OrionBiX are highly sensitive to full parameter updates and prone to overfitting. Collectively, these observations establish that careful alignment between model architecture and adaptation strategy is critical for maximizing the performance of tabular foundation models.
@@
  • OVERALL LEADERBOARD: Across all benchmark suites, the strongest overall performers are TabPFN and OrionMSP. The TabPFN family (including TabPFN-2.0 and TabPFN-2.5) achieves the best aggregate results, with mean accuracies between 0.85–0.88 and weighted F1 scores of 0.84–0.88, reflecting exceptional cross-benchmark consistency. OrionMSP follows closely (overall rank 4.61 in zero-shot and 2.26 in meta-learning), reaching ACC $\approx 0.84$$0.88$ and F1 $\approx 0.83$$0.88$, with particularly strong generalization on OpenML-CC18 and TabZilla.
  • ZERO-SHOT PERFORMANCE: The TabPFN family and OrionMSP are the strongest zero-shot models. TabPFN-2.5 attains the best aggregate zero-shot rank (4.13), with ACC $\approx\textsc{0.85}$0.88 across TALENT, OpenML-CC18, and TabZilla. OrionMSP is a close second (rank 4.61), matching TabPFN’s accuracy on OpenML-CC18 and TabZilla while slightly trailing on TALENT. Both substantially outperform classical baselines by about 2–4 accuracy points.
@@ 6.2.2  Scalability Analysis
Models Small (<1K) Medium (1K-10K) Large (>10K)
Rank ACC F1 Rank ACC F1 Rank ACC F1
@@
XGBoost 8.66 0.8168 0.7964 8.28 0.8363 0.8314 6.52 0.8961 0.8911
CatBoost 8.88 0.8124 0.7935 7.89 0.8340 0.8264 6.43 0.8786 0.8721
Random Forest 9.48 0.7988 0.8187 8.57 0.8285 0.8221 8.68 0.8680 0.8614
LightGBM 8.80 0.8143 0.7789 8.30 0.8314 0.8226 6.82 0.8817 0.8753
TabICL 6.96 0.8301 0.8338 5.86 0.8486 0.8398 5.42 0.8790 0.8731
OrionBiX 7.34 0.8330 0.8150 6.71 0.8348 0.8260 5.54 0.8716 0.8656
OrionMSP 6.76 0.8232 0.8194 4.53 0.8494 0.8402 3.66 0.8831 0.8754
TabPFN-2.0 7.42 0.8325 0.8131 4.63 0.8557 0.8462 6.75 0.8771 0.8701
TabPFN-2.5 5.78 0.8349 0.8194 4.07 0.8566 0.8470 3.42 0.8887 0.8831
Mitra 15.24 0.4334 0.3236 12.89 0.3600 0.2553 12.40 0.3879 0.2801
ContextTab 9.76 0.8179 0.8085 6.82 0.8430 0.8334 9.34 0.8548 0.8464
TabDPT 6.28 0.8333 0.8271 6.58 0.8424 0.8339 6.18 0.8821 0.8754
@@
TabICL 4.68 0.7912 0.7806 3.64 0.8342 0.8256 3.06 0.8431 0.8360
OrionBiX 2.78 0.8149 0.8123 3.99 0.8283 0.8204 4.50 0.8410 0.8306
OrionMSP 3.21 0.8211 0.8151 2.34 0.8413 0.8330 1.38 0.8787 0.8721
TabPFN-2.0 3.31 0.8336 0.8256 2.13 0.8638 0.8548 2.76 0.8748 0.8668
Mitra 6.39 0.6417 0.5988 6.10 0.6083 0.5516 5.82 0.6585 0.6132
TabDPT 3.85 0.8224 0.8188 4.09 0.8289 0.8212 3.54 0.8682 0.8613
@@
TabICL 5.65 0.6746 0.6225 4.86 0.7553 0.7187 5.25 0.6693 0.6212
OrionBiX 5.20 0.7158 0.6792 4.04 0.7718 0.7386 4.03 0.6697 0.6272
OrionMSP 2.90 0.8135 0.7920 2.85 0.8036 0.7798 2.96 0.7231 0.6828
TabPFN-2.0 2.68 0.8216 0.8109 1.93 0.8580 0.8485 1.57 0.8572 0.8459
Mitra 6.48 0.5625 0.4408 6.05 0.5133 0.4116 5.42 0.5720 0.4501
TabDPT 2.66 0.8227 0.8153 2.87 0.8377 0.8284 2.65 0.8120 0.8002
@@
TabICL 4.80 0.7518 0.7397 4.25 0.7251 0.7114 4.29 0.6962 0.6799
OrionBiX 3.33 0.8093 0.8042 2.86 0.7995 0.7939 2.29 0.7848 0.7665
OrionMSP 2.28 0.8275 0.8213 2.10 0.8134 0.8019 2.45 0.8013 0.7817
Mitra 5.21 0.6510 0.6012 4.57 0.6089 0.5599 4.55 0.6304 0.5795
TabDPT 2.92 0.8191 0.8157 2.38 0.8121 0.8049 1.83 0.8407 0.8299
@@
TabICL 3.53 0.7957 0.7710 3.73 0.6424 0.5686 3.71 0.6084 0.5020
OrionBiX 4.38 0.7576 0.7102 3.99 0.6585 0.5892 3.75 0.6327 0.5381
OrionMSP 1.73 0.8298 0.8132 2.06 0.7190 0.6597 2.04 0.6931 0.6152
Mitra 5.53 0.5244 0.4133 4.59 0.5008 0.4013 3.98 0.5459 0.4476
TabDPT 2.85 0.8269 0.8182 1.67 0.8332 0.8245 2.02 0.7686 0.7418
We analyze how model performance varies with dataset size, feature dimensionality, and class imbalance to inform model selection under different data regimes to highlight which models and strategies perform best across varying data constraints.
@@
  • On small datasets (<<1K samples), TabPFN-2.5 achieves the highest zero-shot accuracy (ACC=0.8349, F1=0.8194), marginally ahead of TabPFN-2.0 under meta-learning (0.8336/0.8256) and TabDPT (0.8333/0.8271). OrionMSP (PEFT–Meta, 0.8275/0.8213) follows closely, confirming that transformer backbones still generalize well under constrained data. Fine-tuning TabICL or OrionBiX on small datasets causes sharp degradation (ACC drops >>15%), revealing overfitting risk.
@@
  • On large-scale datasets (>>10K), tree-based ensembles remain strong—XGBoost reaches (0.8911/0.8961)—but transformer TFMs scale competitively, with OrionMSP zero-shot (0.8754/0.8831) and TabDPT zero-shot (0.8821/0.8754). Across all sizes, Mitra consistently underperform (ACC <0.65<0.65), reflecting limited scalability.
@@
Models Narrow (<10) Medium (10-100) Wide (>100)
Rank ACC F1 Rank ACC F1 Rank ACC F1
@@
XGBoost 7.79 0.8199 0.8134 8.27 0.8482 0.8410 6.14 0.9101 0.9039
CatBoost 6.57 0.8124 0.8045 8.26 0.8441 0.8344 6.85 0.9127 0.9084
Random Forest 8.31 0.7981 0.7919 8.82 0.8410 0.8235 9.11 0.8987 0.8936
LightGBM 7.17 0.8107 0.8004 8.27 0.8458 0.8326 9.00 0.8957 0.8908
TabICL 6.01 0.8188 0.8098 5.62 0.8627 0.8549 7.93 0.8948 0.8936
OrionBiX 5.56 0.8089 0.8020 6.70 0.8510 0.8417 8.20 0.8859 0.8849
OrionMSP 3.16 0.8376 0.8295 4.98 0.8572 0.8478 7.00 0.8860 0.8837
TabPFN-2.0 6.08 0.8167 0.8071 4.93 0.8676 0.8589 7.44 0.9129 0.9111
TabPFN-2.5 4.63 0.8243 0.8120 3.92 0.8697 0.8609 4.20 0.9248 0.9240
Mitra 12.52 0.3769 0.2720 13.14 0.3886 0.2781 14.46 0.2521 0.1497
ContextTab 8.64 0.7965 0.7862 7.29 0.8441 0.8452 8.38 0.9008 0.8999
TabDPT 5.35 0.8243 0.8169 6.63 0.8566 0.8483 8.64 0.8845 0.8820
@@
TabICL 3.66 0.7867 0.7779 3.48 0.8383 0.8338 3.86 0.8319 0.8667
OrionBiX 3.92 0.7956 0.7900 3.89 0.8316 0.8278 3.58 0.8429 0.8640
OrionMSP 1.90 0.8253 0.8190 2.21 0.8514 0.8457 2.86 0.8564 0.8737
TabPFN-2.0 2.73 0.8264 0.8171 2.37 0.8683 0.8579 2.28 0.9172 0.9533
Mitra 5.67 0.6646 0.6247 6.12 0.6004 0.5763 6.17 0.4090 0.3224
TabDPT 3.56 0.8129 0.8071 3.82 0.8405 0.8350 3.48 0.8400 0.8497
@@
TabICL 5.28 0.6609 0.6195 5.04 0.7507 0.7101 3.95 0.7757 0.7269
OrionBiX 4.03 0.6922 0.6576 4.33 0.7643 0.7294 3.75 0.7560 0.7046
OrionMSP 2.63 0.7540 0.7297 2.99 0.8025 0.7750 2.90 0.8148 0.7808
TabPFN-2.0 1.86 0.8258 0.8162 2.06 0.8578 0.8469 1.40 0.9346 0.9335
Mitra 5.81 0.5619 0.4478 6.09 0.5349 0.4268 5.60 0.3307 0.2436
TabDPT 2.50 0.8078 0.7948 2.88 0.8378 0.8294 3.40 0.8623 0.8587
@@
TabICL 4.58 0.6375 0.6375 4.21 0.7601 0.7444 4.05 0.7714 0.7622
OrionBiX 2.51 0.7584 0.7584 2.92 0.8147 0.8057 2.50 0.8251 0.8243
OrionMSP 2.18 0.7836 0.7836 2.20 0.8254 0.8107 2.66 0.8293 0.8256
Mitra 4.15 0.6471 0.6471 4.83 0.6177 0.5650 4.61 0.4129 0.3338
TabDPT 2.22 0.7979 0.7979 2.32 0.8299 0.8209 2.94 0.8416 0.8376
@@
TabICL 3.73 0.6196 0.5287 3.74 0.6596 0.5902 2.77 0.8760 0.8502
OrionBiX 3.81 0.6278 0.5444 4.08 0.6851 0.6177 4.00 0.6737 0.6166
OrionMSP 2.10 0.6795 0.6113 1.97 0.7477 0.6939 1.88 0.8042 0.7663
Mitra 4.28 0.5518 0.4481 4.78 0.5106 0.4108 4.77 0.2831 0.1874
TabDPT 1.72 0.7821 0.7631 2.02 0.8305 0.8201 1.55 0.8925 0.8901
@@
  • For narrow datasets (<<10 features), performance converges across models: OrionMSP zero-shot (0.8376/0.8295) slightly leads, followed by TabDPT and TabPFN, while classical baselines remain competitive.
  • On medium-width data (10–100 features), TabPFN dominates under both zero-shot (0.8376/0.8589) and meta-learning (0.8683/0.8579), outperforming all baselines and transformer variants.
  • The advantage of TFMs becomes most pronounced on wide feature spaces (>>100 features): TabPFN-2.0 achieves exceptional results under SFT (ACC=0.9346, F1=0.9335) and meta-learning (0.9172/0.9533), far exceeding other models. TabICL and OrionBiX degrade when fully fine-tuned, confirming that dense feature spaces accentuate overfitting. Traditional models (XGBoost) remain competitive but trail by 2–3 %.
@@
XGBoost 8.44 0.8175 0.8110 7.28 0.8853 0.8779
CatBoost 8.60 0.8076 0.8020 6.69 0.8780 0.8669
Random Forest 9.36 0.7983 0.7955 7.99 0.8736 0.8639
LightGBM 8.70 0.8071 0.7977 7.31 0.8769 0.8626
TabICL 5.81 0.8279 0.8233 6.00 0.8801 0.8692
OrionBiX 6.95 0.8096 0.8040 6.07 0.8782 0.8677
OrionMSP 5.20 0.8265 0.8202 3.96 0.8835 0.8725
TabPFN-2.0 4.73 0.8367 0.8309 6.23 0.8803 0.8691
TabPFN-2.5 3.81 0.8406 0.8348 4.49 0.8835 0.8727
Mitra 13.61 0.2763 0.1540 12.48 0.4824 0.3891
ContextTab 7.33 0.8152 0.8095 8.19 0.8722 0.8601
TabDPT 6.27 0.8233 0.8189 6.65 0.8793 0.8684
@@
Discussion : Scalability analysis underscores that TabPFN, OrionMSP, and TabDPT each occupy distinct strengths. TabPFN excels on small, medium, and high-dimensional data where Bayesian priors and strong supervision prevail; OrionMSP scales best with large datasets and imbalanced distributions due to meta-learning robustness; and TabDPT sustains near-top performance across all settings with minimal computational overhead. Classical baselines remain reliable but consistently trail the best TFMs by 2–4% in accuracy.
@@
  • DATASET SIZE: Small datasets favor the TabPFN family (TabPFN-2.5 zero-shot, TabPFN-2.0 meta-learning), with TabDPT as a close, efficient alternative; medium datasets favor TabPFN (meta & SFT); large datasets see XGBoost slightly leading in raw accuracy, with OrionMSP and TabDPT close behind.
@@
  • Class imbalance: Meta-learning fine-tuning yields the best robustness under imbalance, with OrionMSP (Rank-$2.16$, ACC = $0.8735$, F1 = $0.8636$) outperforming other TFMs; TabPFN (rank $2.50$) remains second.